OpenStudy (anonymous):

Would someone help, and not throw a rock at meXD Write the following equation in Standard Form: 9y = -6x + 14 A. -6x + 9y = 14 B. 6x + 9y = 14 C. -6x + 9y = 7 D. 6x - 9y = 14

OpenStudy (vanessaperezxo):

Ok, i can help.

OpenStudy (vanessaperezxo):

Which one do you think is the answer?

OpenStudy (anonymous):

-6x + 9y = 14

OpenStudy (anonymous):

@vanessaperezxo

OpenStudy (sleepyhead314):

hmm to get the x to the other side you should try to add 6x to both sides 9y = -6x + 14 +6x +6x

OpenStudy (anonymous):

15xy = 0 + 14.-.?

OpenStudy (vanessaperezxo):

Ok, what you need to do next is move the -6x to the other side. 

OpenStudy (sleepyhead314):

uh oh! be careful, you can't add things that have different variables

OpenStudy (vanessaperezxo):

Sorry the bottom number is 14 not 16

OpenStudy (vanessaperezxo):

do any of the answers look similar to this?

OpenStudy (anonymous):

B.

OpenStudy (vanessaperezxo):

Yes, your correct. Do you understand why its B?

OpenStudy (anonymous):

6x + 9y = 14? is the answerr

OpenStudy (sleepyhead314):

yes

OpenStudy (anonymous):

yes